One number is 10 more than the other number. Twice the smaller number is 13 more than the larger number. What are the numbers?
Answer by Paul(988)

Smaller nukber =x
Larger number = 10+x
2(x)=13+10+x
2x=23+x
x=23
23+10=33

hence the numbers are 23 and 33.
Paul.